Here are data on two stocks, both of which have discount rates of 12%:

                                            Stock A     Stock B
ROE                                       12 %           10 %
Earnings per share              $ 2.00         $ 1.70
Dividends per share            $ 1.20         $ 1.20
________________________________________

a. What are the dividend "payout" and "plowback" ratios for each firm? 

b. What are the expected dividend growth rates for each firm?

c. What is the proper stock price for each firm? 
Answer #1

a). Dividend payout Ratio = Dividend per share/Earning per share

Plowback ratio = 1- Dividend payout Ratio

Stock A

Dividend payout Ratio = $1.20/$2.00 = 0.60 or 60%

Plowback ratio = 1- 0.60 = 0.40 or 40%

Stock B

Dividend payout Ratio = $1.20/$1.7 = 0.7059 or 70.59%

Plowback ratio = 1- 0.7059 = 0.2941 or 29.41%

b). Growth rate = ROE*plowback ratio

Stock A

Growth rate = 12%*40% = 4.8%

Stock B

Growth rate = 10%*29.41% = 2.941%

c). Discount rate (ke)= 12%

Current Price = D0(1+g)/Ke-g

Stock A

D0 = $1.2

g = Growth rate = 4.8%

Current Price = 1.2(1+0.048)/(0.12-0.048)

= $ 17.467 per share

Stock B

D0 = $1.2

g = Growth rate = 2.941%

Current Price = 1.2(1+0.02941)/(0.12-0.02941

= $ 13.636 per share

If you need any clarification, you can ask in comments.

If you like my answer, then please up-vote as it will be motivating